OpenStudy (anonymous):

a ball is hit into the air; its height, in metres. after t seconds is h = -5^2 + 30t + 1 for how long is the ball in the air? i understand i must find the x intercepts, but how do i do so by factoring (without quadratic formula) with a smaller c than b?

OpenStudy (anonymous):

is that -5t^2 as the first term?

OpenStudy (turingtest):

assuming it is it is unfactorable by traditional means http://www.wolframalpha.com/input/?i=+factor+-5t%5E2+%2B+30t+%2B+1

OpenStudy (turingtest):

by traditional I mean high school

OpenStudy (anonymous):

yes -5t^2 is the first term, and yes in high school terms i don't get how i'm supposed to get the answer? if i can't factor it (i assume i can't in this case) what do i do?

OpenStudy (turingtest):

You know, we have a moderator on here who gave me a way to do it, but I can't find the file right now. The point is that in high-school terms that practically can't be done. I'm not sure why you would be asked to answer something like that without prior knowledge of the technique involved.

OpenStudy (agreene):

This is definitely a high school (or lower) level completing the square question. divide off the -5 and we have t^2-6t-1/5=0 add the 1/5 over, t^2-6t=1/5 Now we need constant of 9 to factor, so we add it to both sides: t^2-6t+9 = 46/5 (t-3)^2 = 46/5 ... etc.

OpenStudy (turingtest):

I always hated completing the square, I don't know why. I was only focused on factoring. Agreene's way is probably what your class is looking for.